# Data Augmentation for Traffic Classification

### 19 Jan 2024 v1

@@ 3.4 Combining augmentations (G4)
To address G1, each trained model is associated to an individual augmentation. However, in CV it is very common to combine multiple augmentations [8]. Hence, we aim to complement G1 by measuring the performance of three different policies augmenting mini-batches based on a set $\mathcal{A^{\prime}}\subset\mathcal{A}$ composed of top-performing augmentations based on G1 benchmark: the Ensemble policy independently and uniformly samples one of the augmentations in $\mathcal{A^{\prime}}$ to apply to each sample in a mini-batch; the RandomStack policy randomly samples a different order for applying all $\mathcal{A^{\prime}}$ augmentations to each sample in a mini-batch; finally, the MaskedStack policy uses a predefined order for applying $\mathcal{A^{\prime}}$ augmentations but each augmentation has its own probability to be applied.
@@ 4.1 Datasets
$\mathtt{MIRAGE}$-$19$ [3] is a public dataset gathering traffic logs from $20$ popular Android apps [^2] Despite being advertised of having traffic from 40 apps, the public version of the dataset only contains 20 apps. as collected by the ARCLAB laboratories of the University of Napoli Federico II. Multiple measurement campaigns were operated by instrumenting $3$ Android devices handed off to $\approx$300300300300 volunteers (students and researchers) for interacting with the selected apps for short sessions. Each session resulted in a pcap file and an strace log mapping each socket to the corresponding Android application name. Pcaps were then post-processed to obtain bidirectional flow logs by grouping all packets belonging to the same 5-tuple (srcIP, srcPort, dstIP, dstPort, L4proto) and extracting both aggregate metrics (e.g., total bytes, packets, etc.), per-packet time series (packet size, direction, TCP flags, etc.), raw packets payload bytes (encoded as a list of integer values) and mapping a ground-truth label by means of the strace logs.
@@
$\mathtt{Enterprise}$ is instead a private [^3] Due to NDA we are not allowed to share the dataset. dataset collected by monitoring network flows from vantage points deployed in residential access and enterprise campus networks. For each flow, the logs report multiple aggregate metrics (number of bytes, packets, TCP flags counters, round trip time statistics, etc.), and the packet time series of packet size, direction and IAT for the first 50 packets of the flow. Each flow record is also enriched with an application label provided by a commercial DPI software directly integrated into the monitoring solution and supporting hundreds of applications and services.
@@
Data folds and normalization. As described in Sec. 3.1, each flow is modeled via a multivariate time series $\rm{\bf x}$ considering $D=3$ features (packets size, direction, and IAT) related to the first $T=20$ packets (applying zero padding in the tail where needed). From the curated datasets we created 80 random 70/15/15 train/validation/test folds. We then processed each train+val split to extract statistics that we used for normalizing the data and to drive the augmentation process. Specifically, we computed both per-coordinate $(d,t)$ and global (i.e., flattening all flows time series into a single array) mean and standard deviation for each class—these statistics provided us the $\sigma_{(d,t)}^{y}$ and $\sigma_{(:,t)}^{y}$ needed for the augmentations (cf. Fig. 1 and Table 1-3). For IAT, we also computed the global 99th percentile across all classes $q_{iat}^{99}$. Given a multi-variate input $\rm{\bf x}$, we first clip packet size values in the range $[0,1460]$ and IAT values in the range [1e-7, $q_{iat}^{99}$]. Due to high skew of IAT distributions, we log10-scaled the IAT feature values. [^4] We do not log-scale packet sizes values as this can reduce accuracy based on preliminary empirical assessments. Last, all features are standardized to provide values $x_{(d,t)}\in[0,1]$.
Model architecture and training. We rely on a 1d-CNN based neural network architecture with a backbone including 2 ResNet blocks followed by a linear head resulting in a quite compact architecture of $\approx$100k parameters. (see Fig. 7 and Listing 1 in the appendix for details). Models are trained for a maximum of 500 epochs with a batch size B=1,024 via an AdamW optimizer with a weight decay of 0.0001 and a cosine annealing learning rate scheduler initialized at 0.001. Training is subject to early stopping by monitoring if validation accuracy does not improve within 20 epochs by 0.02. We coded our modeling framework using PyTorch and PyTorch Lightning and ran our modeling campaigns on Linux servers equipped with multiple NVIDIA Tesla V100 GPUs. We measure classification performance via the weighted F1 score considering a reference baseline where training is not subject to augmentations.
@@ 5.1 Augmentations benchmark (G1)
We start by presenting the overall performance of the selected augmentations. Specifically, Table 5 collects results obtained by applying augmentation via Inject with $N_{inject}=1$ (i.e., each original sample is augmented once) [^5] Since we train the reference baseline with a batch size B=1024, when adding augmentations we instead adopt B=512 (which doubles via injection). and sampling uniformly the magnitude $\alpha\sim U[0,1]$. Values reported correspond to the average weighted F1 score across 80 runs and related 95th-percentile confidence intervals.
Table 5: Augmentations benchmark.
@@
Figure 3: Augmentations rank and critical distance.
@@
Augmentations rank. Overall, all augmentations are beneficial but Horizontal Flip which, as we shall see in Sec. 5.3, breaks class semantics. As expected, not all augmentations provide the same gain and their effectiveness may vary across datasets. Specifically, sequence and masking better suit our TC tasks.
For a more fine-grained performance comparison, we complement Table 5 results by analyzing augmentations rank via a critical distance by following the procedure described in [12]. Specifically, for each of the 80 modeling runs we first ranked the augmentations from best to worst (e.g., if augmentations A, B, and C yield a weighted F1 of 0.9, 0.7, and 0.8, their associated rankings would be 1, 3, and 2) splitting ties using the average ranking of the group (e.g., if augmentations A, B, and C yield a weighted F1 of 0.9, 0.9 and 0.8, their associated rankings would be 1.5, 1.5, and 3). This process is then repeated across the 80 runs and a global rank is obtained by computing the mean rank for each augmentation. Last, these averages are compared pairwise using a post-hoc Nemenyi test to identify which groups of augmentations are statistically equivalent. This decision is made using a Critical Distance $CD=q_{\alpha}\sqrt{k(k+1)/6N}$, where $q_{\alpha}$ is based on the Studentized range statistic divided by $\sqrt{2}$, $k$ is equal to the number of augmentations compared and $N$ is equal to the number of samples used. Results are then collected in Fig. 3 where each augmentation is highlighted with its average rank (the lower the better) and horizontal bars connect augmentations that are statistically equivalent. For instance, while Table 5 shows that Translate is the best on average, Fig. 3 shows that {Translate, Wrap, Permutation, Dup-FastRetr} are statistically equivalent. We remark that CD is reported for $\mathtt{MIRAGE}$-$19$ and $\mathtt{MIRAGE}$-$22$ but similar considerations hold for $\mathtt{Enterprise}$.
As previously reported, our training process is subject to an early stop mechanism. Interestingly, we observed that augmentations yielding better performance also present a longer number of training epochs (see Fig. 8 in Appendix). This suggests that effective augmentations foster the extraction of better data representations, but some studies in CV also show that early stopping might not necessarily be the best option to achieve high accuracy in some scenarios. An in-depth investigation of these training mechanisms is however outside the scope of this paper.
@@
Sensitivity to magnitude. Most of the augmentations we analyzed are subject to a magnitude $\alpha$ hyper-parameter (see Table 1-3) that is randomly selected for the results in Table 5. To investigate the relationship between classification performance and augmentation magnitude we selected 3 augmentations among the top performing ones {Translation, Wrap, Permutation} and three among the worst performing {Gaussian Noise, Sine WrapUp, Constant WrapUp}. [^7] We exclude HorizontalFlip as it hurts performance and Interpolation since it does not depend on a magnitude. For each augmentation, we performed 10 modeling runs using magnitude $\alpha=0.5$ and we contrasted these results with the related runs from the previous modeling campaign. Specifically, by grouping all results we obtained a binary random-vs-static performance comparison which we investigated through a Wilcoxon signed rank sum test that indicated no statistical difference, i.e., the selection of magnitude is not a distinctive factor to drive the augmentation performance. The same conclusion holds true when repeating the analysis for each individual augmentation rather than grouping them together.
Takeaways. Although we do not observe any dependency on the augmentation magnitude $\alpha$ (as will be discussed further in Sec. 5.3), augmentations performance can still be affected by their tuning. Unfortunately, this tuning process often relies on trial-and-error, making it challenging to operate manually.
@@ 5.2 Training batches composition (G2)
Figure 4: Comparing Replace, Inject and Pre-augment batch creation policies.
@@
Table 6: Impact of class-weighted sampler on $\mathtt{MIRAGE}$-$19$.
@@
Takeaways. Paying too much attention to minority classes can perturb the overall classifier balance, so we discourage training with a class-weighted sampler.
@@
Figure 5: Investigating the relationship between train, augmented and test samples.
@@ 5.3 Latent space geometry (G3)
Augmented-vs-Test samples. We start our analysis by taking the point of view of the test samples. Specifically, we investigated which type of points are found in the “neighbourhood” of a test sample. To do so, we started creating “true anchors” by projecting both the original training data and 5 augmentations of each training sample—these anchors are “proxy” of what is presented to the model during training. Then we projected the test samples and looked for the closest 10 anchors (based on cosine similarity) of each test sample. Finally, we counted how many of the 10 anchors share the same label as the test samples. Results for each augmentation are reported in Fig. 5 for $\mathtt{MIRAGE}$-$19$ and $\mathtt{MIRAGE}$-$22$ (similar results holds for $\mathtt{Enterprise}$ dataset) as a scatter plot where the coordinates of each point correspond to the average number of anchors with the correct label found and their average cosine similarity with respect to the test sample. Each augmentation is color-coded with respect to its weighted F1 score.
@@
Figure 6: Comparing original and augmented samples in the latent space.
@@
Last, for each test sample we looked at the closest augmented anchor and the closest original sample anchor with the same label. The average ratio of those pairwise distances is centered around 1—augmented samples proxy test samples as much as the original samples.
@@
Takeaways. Augmentations with higher performance do not better mimic test samples compared with original samples. Rather, they help training the feature extractor f()normal-$f(\cdot)$ so that projected test samples are found in neighborhood of points likely to have the expected label.
@@
Augmented-vs-Original samples. We complement the previous analysis by investigating original $\bf x$ and augmented $\bf x^{\prime}$ samples relationships. Differently from before, for this analysis original samples are augmented once. Then all points are projected in the latent space $f({\bf x})$ and $f({\bf x^{\prime}})$ and visualized by means of a 2d t-SNE projection.111111 Our model architecture uses a latent space of 256 dimensions (see Listing 1) which the t-SNE representation compresses into a 2d space. We also compute the Kernel Density Estimation (KDE) of the Euclidean distance across all pairs. Fig. 6 presents the results for 2 top-performing (Translate, Wrap) and 4 poor-performing (Constant Wrapup, Interpolation, Sine Wrapup, Horizontal Flip) augmentations for $\mathtt{MIRAGE}$-$19$ dataset. Points in the t-SNE charts are plotted with alpha transparency, hence color saturation highlights prevalence of either augmented or original samples.
@@
Linking back to the previous observations about Horizontal Flip and Interpolation, results now show the more “aggressive” nature of Interpolation—the t-SNE chart is split vertically with the left (right) side occupied by augmented (original) samples only and Euclidean distance KDEs show heavier tails. By recalling their definition, while it might be easy to realize why Horizontal Flip is a poor choice—a client will never observe the end of the flow before seeing the beginning—it is difficult to assess a priori the effect of Interpolation. Overall, both augmentations break class semantics.
@@
At the opposite side of the performance range, we find augmentations like Sine WrapUp and Constant WrapUp. From Fig. 6 we can see that both introduce little-to-no variety—the euclidean distance distributions are centered around zero. That said, comparing their t-SNE charts we can still observe a major difference between the two transformations which relates to their design. Specifically, Constant WrapUp is applied only to IAT and introduces negligible modifications to the original samples. Conversely, Sine WrapUp is applied on either packet size or IAT. As for Constant WrapUP, when acting on IAT changes are subtle, while variations of size lead to generating an extra “mode” (notice the saturated cluster of points on the left side of the t-SNE plot). In other words, besides the design of the augmentation itself, identifying a good parametrization is very challenging and in this case is also feature-dependent.
@@
Table 7: Combining augmentations.
@@
Takeaways. Effective transformations operate in a “sweet spot”: they neither introduce too little variety—traditional policies like Random Over Sampling (ROS) and Random Under Sampling (RUS) [23] are ineffective—nor break classes semantic by introducing artificial “modes”.

### 23 Jan 2024 v2

@@ 3.4 Combining augmentations (G4)
To address G1, each trained model is associated to an individual augmentation. However, in CV it is very common to combine multiple augmentations [8]. Hence, we aim to complement G1 by measuring the performance of three different policies augmenting mini-batches based on a set $\mathcal{A^{\prime}}\subset\mathcal{A}$ composed of top-performing augmentations based on the G1 benchmark: the Ensemble policy uniformly samples one of the augmentations in $\mathcal{A^{\prime}}$ independently for each mini-batch sample; the RandomStack policy randomly shuffles $\mathcal{A^{\prime}}$ independently for each mini-batch sample before applying all augmentations; finally, the MaskedStack policy uses a predefined order for $\mathcal{A^{\prime}}$ but each augmentation is associated to a masking probability, i.e., each sample in the mini-batch independently selects a subset of augmentations of the predefined order.
@@ 4.1 Datasets
$\mathtt{MIRAGE}$-$19$ [3] is a public dataset gathering traffic logs from $20$ popular Android apps [^2] Despite being advertised with having traffic from 40 apps, the public version of the dataset only contains 20 apps. collected at the ARCLAB laboratory of the University of Napoli Federico II. Multiple measurement campaigns were operated by instrumenting $3$ Android devices handed off to $\approx$300300300300 volunteers (students and researchers) for interacting with the selected apps for short sessions. Each session resulted in a pcap file and an strace log mapping each socket to the corresponding Android application name. Pcaps were then post-processed to obtain bidirectional flow logs by grouping all packets belonging to the same 5-tuple (srcIP, srcPort, dstIP, dstPort, L4proto) and extracting both aggregate metrics (e.g., total bytes, packets, etc.), per-packet time series (packet size, direction, TCP flags, etc.), raw packets payload bytes (encoded as a list of integer values) and mapping a ground-truth label by means of the strace logs.
@@
$\mathtt{Enterprise}$ is instead a private [^3] Due to NDA we are not allowed to share the dataset. dataset collected by monitoring network flows from vantage points deployed in residential access and enterprise campus networks. For each flow, the logs report multiple aggregate metrics (number of bytes, packets, TCP flags counters, round trip time statistics, etc.), and the packet time series of packet size, direction and IAT for the first 50 packets of each flow. Moreover, each flow record is also enriched with an application label provided by a commercial DPI software directly integrated into the monitoring solution and supporting hundreds of applications and services.
@@
Data folds and normalization. As described in Sec. 3.1, each flow is modeled via a multivariate time series $\rm{\bf x}$ consisting of $D=3$ features (packets size, direction, and IAT) related to the first $T=20$ packets (applying zero padding in the tail where needed). From the curated datasets we created 80 random 70/15/15 train/validation/test folds. We then processed each train+val split to extract statistics that we used for normalizing the data and to drive the augmentation process. Specifically, we computed both per-coordinate $(d,t)$ and global (i.e., flattening all flows time series into a single array) mean and standard deviation for each class—these statistics provided us the $\sigma_{(d,t)}^{y}$ and $\sigma_{(:,t)}^{y}$ needed for the augmentations (see Fig. 1 and Tables 1-3). For IAT, we also computed the global 99th percentile across all classes $q_{iat}^{99}$. Given a multi-variate input $\rm{\bf x}$, we first clip packet size values in the range $[0,1460]$ and IAT values in the range [1e-7, $q_{iat}^{99}$]. Due to high skew of IAT distributions, we also log10-scaled the IAT feature values. [^4] We did not log-scale packet sizes values as we found this can reduce accuracy based on preliminary empirical assessments. Last, all features are standardized to provide values $x_{(d,t)}\in[0,1]$.
Model architecture and training. We rely on a 1d-CNN based neural network architecture with a backbone including 2 ResNet blocks followed by a linear head resulting in a compact architecture of $\approx$100k parameters. (see Fig. 7 and Listing 1 in the appendix for details). Models are trained for a maximum of 500 epochs with a batch size B=1,024 via an AdamW optimizer with a weight decay of 0.0001 and a cosine annealing learning rate scheduler initialized at 0.001. Training is subject to early stopping by monitoring if the validation accuracy does not improve by 0.02 within 20 epochs. We coded our modeling framework using PyTorch and PyTorch Lightning and ran our modeling campaigns on Linux servers equipped with multiple NVIDIA Tesla V100 GPUs. We measured the classification performance via the weighted F1 score considering a reference baseline where training is not subject to augmentations.
@@ 5.1 Augmentations benchmark (G1)
We start by presenting the overall performance of the selected augmentations. Specifically, Table 5 collects results obtained by applying augmentations via Inject with $N_{inject}=1$ (i.e., each original sample is augmented once) [^5] Since we train the reference baseline with a batch size B=1024, when adding augmentations we instead adopt B=512 (which doubles via injection). and sampling uniformly the magnitude $\alpha\sim U[0,1]$. Table 5 shows the average weighted F1 score across 80 runs and related 95th-percentile confidence intervals.
Table 5: Augmentations benchmark (G1).
@@
Figure 3: Augmentations rank and critical distance (G1).
@@
Augmentations rank. Overall, all augmentations are beneficial except for Horizontal Flip which, as we shall see in Sec. 5.3, breaks class semantics. As expected, not all augmentations provide the same gain and their effectiveness may vary across datasets. Specifically, sequence and masking better suit our TC tasks.
For a more fine-grained performance comparison, we complement Table 5 results by analyzing augmentations rank via a critical distance by following the procedure described in [12]. Specifically, for each of the 80 modeling runs we first ranked the augmentations from best to worst (e.g., if augmentations A, B, and C yield a weighted F1 of 0.9, 0.7, and 0.8, their associated rankings would be 1, 3, and 2) splitting ties using the average ranking of the group (e.g., if augmentations A, B, and C yield a weighted F1 of 0.9, 0.9 and 0.8, their associated rankings would be 1.5, 1.5, and 3). This process is then repeated across the 80 runs and a global rank is obtained by computing the mean rank for each augmentation. Last, these averages are compared pairwise using a post-hoc Nemenyi test to identify which groups of augmentations are statistically equivalent. This decision is made using a Critical Distance $CD=q_{\alpha}\sqrt{k(k+1)/6N}$, where $q_{\alpha}$ is based on the Studentized range statistic divided by $\sqrt{2}$, $k$ is equal to the number of augmentations compared and $N$ is equal to the number of samples used. Results are then collected in Fig. 3 where each augmentation is highlighted with its average rank (the lower the better) and horizontal bars connect augmentations that are statistically equivalent. For instance, while Table 5 shows that Translate is the best on average, Fig. 3 shows that {Translate, Wrap, Permutation, Dup-FastRetr} are statistically equivalent. We remark that Fig. 3 refers to $\mathtt{MIRAGE}$-$19$ and $\mathtt{MIRAGE}$-$22$ but similar considerations hold for $\mathtt{Enterprise}$ as well.
Recall that our training process is subject to an early stop mechanism. Interestingly, we observed that augmentations yielding better performance also present a longer number of training epochs (see Fig. 8 in Appendix). This hints that effective augmentations foster better data representations extraction, although some CV studies also show that early stopping might not necessarily be the best option to achieve high accuracy in some scenarios. An in-depth investigation of these training mechanisms is however out of scope for this paper.
@@
Sensitivity to magnitude. Most of the augmentations we analyzed are subject to a magnitude $\alpha$ hyper-parameter (see Tables 1-3) that is randomly selected for the results in Table 5. To investigate the relationship between classification performance and augmentation magnitude we selected 3 augmentations among the top performing ones {Translation, Wrap, Permutation} and three among the worst performing {Gaussian Noise, Sine WrapUp, Constant WrapUp}. [^7] We excluded HorizontalFlip as it hurt performance and Interpolation since it does not depend from a magnitude. For each augmentation, we performed 10 modeling runs using magnitude $\alpha=0.5$ and we contrasted these results with the related runs from the previous modeling campaign. Specifically, by grouping all results we obtained a binary random-vs-static performance comparison which we investigated through a Wilcoxon signed rank sum test that indicated no statistical difference, i.e., the selection of magnitude is not a distinctive factor to drive the augmentation performance. The same conclusion holds true when repeating the analysis for each individual augmentation rather than grouping them together.
Takeaways. Although we do not observe any dependency on the augmentation magnitude $\alpha$, augmentations performance can still be affected by their tuning (as will be discussed further in Sec. 5.3). Unfortunately, this tuning process often relies on a trial-and-error process, making it challenging to operate manually.
@@ 5.2 Training batches composition (G2)
Figure 4: Comparing Replace, Inject and Pre-augment batch creation policies (G2).
@@
Table 6: Impact of class-weighted sampler on $\mathtt{MIRAGE}$-$19$ (G2).
@@
Takeaways. Paying too much attention to minority classes can perturb the overall classifier balance, so we discourage the use of class-weighted samplers.
@@
Figure 5: Investigating train, augmented and test samples relationships (G3).
@@ 5.3 Latent space geometry (G3)
Augmented-vs-Test samples. We start our analysis by taking the point of view of the test samples. Specifically, we investigated which type of points are found in the “neighbourhood” of a test sample. To do so, we started creating “true anchors” by projecting both the original training data and 5 augmentations of each training sample—these anchors are “proxy” of what is presented to the model during training. Then we projected the test samples and looked for the closest 10 anchors (based on cosine similarity) of each test sample. Finally, we counted how many of the 10 anchors share the same label as the test samples. Results for each augmentation are reported in Fig. 5 for $\mathtt{MIRAGE}$-$19$ and $\mathtt{MIRAGE}$-$22$ (similar results holds for $\mathtt{Enterprise}$) as a scatter plot where the coordinates of each point correspond to the average number of anchors with the correct label found and their average cosine similarity with respect to the test sample. Each augmentation is color-coded with respect to its weighted F1 score.
@@
Figure 6: Comparing original and augmented samples in the latent space (G3).
@@
Last, for each test sample we looked at the closest augmented anchor and the closest original sample anchor with the same label. The average ratio of those pairwise distances is centered around 1—augmented samples “mimic” test samples as much as the original samples do.
@@
Takeaways. Top-performing augmentations do not better mimic test samples compared to original samples. Rather, they help training the feature extractor f()normal-$f(\cdot)$ so that projected test samples are found in neighborhood of points likely to have the expected label.
@@
Augmented-vs-Original samples. We complement the previous analysis by investigating original $\bf x$ and augmented $\bf x^{\prime}$ samples relationships. Differently from before, for this analysis original samples are augmented once. Then all points are projected in the latent space $f({\bf x})$ and $f({\bf x^{\prime}})$ and visualized by means of a 2d t-SNE projection.111111 Our model architecture uses a latent space of 256 dimensions (see Listing 1) which the t-SNE representation compresses into a 2d space. We also compute the Kernel Density Estimation (KDE) of the Euclidean distance across all pairs. Figure 6 presents the results for 2 top-performing (Translate, Wrap) and 4 poor-performing (Constant Wrapup, Interpolation, Sine Wrapup, Horizontal Flip) augmentations for $\mathtt{MIRAGE}$-$19$. Points in the t-SNE charts are plotted with alpha transparency, hence color saturation highlights prevalence of either augmented or original samples.
@@
Linking back to the previous observations about Horizontal Flip and Interpolation, results now show the more “aggressive” nature of Interpolation—the t-SNE chart is split vertically with the left (right) side occupied by augmented (original) samples only and the Euclidean distance KDEs show heavier tails. By recalling their definition, while it might be easy to realize why Horizontal Flip is a poor choice—a client will never observe the end of the flow before seeing the beginning, hence they are too artificious—it is difficult to assess a priori the effect of Interpolation. Overall, both augmentations break class semantics.
@@
At the opposite side of the performance range we find augmentations like Sine WrapUp and Constant WrapUp. From Fig. 6 we can see that both introduce little-to-no variety—the Euclidean distance distributions are centered around zero. That said, comparing their t-SNE charts we can still observe a major difference between the two transformations which relates to their design. Specifically, Constant WrapUp is applied only to IAT and introduces negligible modifications to the original samples. Conversely, Sine WrapUp is applied on either packet size or IAT. As for Constant WrapUP, the changes to IAT are subtle, while variations of packet size lead to generating an extra “mode” (notice the saturated cluster of points on the left side of the t-SNE plot). In other words, besides the design of the augmentation itself, identifying a good parametrization is very challenging and in this case is also feature-dependent.
@@
Table 7: Combining augmentations (G4).
@@
Takeaways. Effective transformations operate in a “sweet spot”: they neither introduce too little variety—traditional policies like Random Over Sampling (ROS) and Random Under Sampling (RUS) [23] are ineffective—nor they break classes semantic by introducing artificial “modes”.